OpenStudy (anonymous):

Someone helps me please. I am blank now. :) If we divide [1/n, 1] into parts with the same length 1/n^2, how many parts can we have?

OpenStudy (anonymous):

I know if we divide [0,1] into n parts, then, each part will have 1/ n length. But temporarily being blank with this.

OpenStudy (anonymous):

$$(1-1/n)\div 1/n^2=n^2(1-1/n)=n(n-1)$$